OpenStudy (anonymous):

Determine the slope between the points (-1, -5) and (5, 4) A. 2/3 B. -1/4 C. 1/4 D. 3/2

OpenStudy (jhannybean):

slope follows the formula \[\large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[(x_{1},y_{1}) \implies (-1,-5)\]\[(x_{2},y_{2}) \implies (5,4)\] can you plug the points into the formula and figure out the slope?

OpenStudy (anonymous):

the calculator i use won't allow subscript, what should i do then?

OpenStudy (jhannybean):

no no, not subscripts, plug in the points xD the arrow means the x1,y1,x2,y2 are represented by those points

OpenStudy (jhannybean):

\[\large \frac{y_{2}-y_{1}}{x_{2}-x_{1}}= \frac{4-(-5)}{5-(-1)}\]

OpenStudy (jhannybean):

Are you able to simplify that now? :)

OpenStudy (anonymous):

ohp yess! sso 3/2 !? :D

OpenStudy (anonymous):

for some odd reason i thought i needed to have the subscripts haha, now i feel ridiculous. but seriously, thank you for the help!! now my homework actually makes sense (:

OpenStudy (jhannybean):

yes good job :)